OpenStudy (anonymous):

Write the steps needed to solve a system by elimination. System: 3x + 2y = 6 4x - y = 8

OpenStudy (ajprincess):

1.Multiply the second equation by 2. 2.Add both the equations.

OpenStudy (anonymous):

3x+2y=6 4x_y=8.....*2

OpenStudy (anonymous):

3x+2y=6 8x-2y=16 so now find the value of x..

OpenStudy (anonymous):

11x=22

OpenStudy (anonymous):

now x value is x=2

OpenStudy (anonymous):

so now we have got the value of x value right